# What Do You See? Enhancing Zero-Shot Image Classification with Multimodal Large Language Models

### v2

@@ 3.2 Ablation Studies
We conducted a series of ablation studies to explore different versions of our method and investigate the impact of each feature, different fusion approaches, and different cross-modal embedding models. Table 2 presents the results of our method using solely the encoded image description, referred to as the description feature (DF), as our input. We also report the results obtained by using encoded initial predictions, termed as the prediction feature (PF), as our input, as well as using both DF and PF concurrently as inputs. Additionally, Table 2 shows the results of employing image feature (IF) alongside DF or PF as inputs, and finally, we present the results when leveraging all available inputs – specifically, DF, PF, and IF.
From the results in Table 2, it is clear that incorporating all three features (DF, PF, and IF) yields the best performance across most datasets, except for the CIFAR datasets (Krizhevsky et al. (2009)). This discrepancy may arise from the low resolution of CIFAR images (originally 3232), where utilizing the IF may degrade accuracy compared to using only DF and PF.
CLIP model (Radford et al. (2021)) | |||
@@
Table 3 shows the results of our second set of ablation studies, where we report the results for 5,000 randomly selected images from the ImageNet dataset (Deng et al. (2009)). We explore the use of different cross-modal embedding models (CLIP [ViT-L/14], CLIP [ViT-B/32], and CLIP [ViT-B/16]) (Radford et al. (2021)), and additionally investigate different fusion approaches. Rather than using the mean feature vector of our input features (DF, PF, and IF), we calculated the similarity between each input feature separately and the dataset class label features. Subsequently, we fused the similarity scores to generate a single similarity score for each class label in each dataset. We explored two fusion methods: averaging and taking the maximum for each class label. As shown, averaging our three features (IF, DF, PF) yields the best results.
4 Computation resources
@@ 5 Limitations
Our method introduces a new approach by leveraging multimodal LLMs to enhance the accuracy of zero-shot image classification. However, it is important to acknowledge that there are still some limitations inherent in our proposed method. Since our method relies on multiple queries to a multimodal LLM to generate the required features (i.e., DF and PF), there may be potential constraints when running on devices with limited computational power, and it may consume more time compared to other methods. Nevertheless, we believe that advancements in LLMs will lead to models that can run efficiently on lower computational power. This would enable broader accessibility and applicability of such models, such as Gemini (Gemini Team Google (2023)), GPT (Brown et al. (2020)), and LLaMA (Touvron et al. (2023)).
@@ 6 Conclusion
In this work, we introduced a zero-shot image classification method that relies on multimodal large language models (LLMs). Our approach involves using a multimodal LLM to describe the input testing image and make an initial class prediction based on input testing image and the target class label names. Subsequently, we fuse the encoded features of the image description, initial LLM’s prediction, and input testing image to retrieve similar encoded features to class labels from the target dataset. Our method is straightforward and easy to implement, resulting in significant improvements in zero-shot classification accuracy when compared with prior methods in this domain.
@@ Appendix A Additional Details
In the main paper, we presented our method for zero-shot image classification. The inference process of our method is concisely described in Algorithm 1. As part of our method, we employed a set of prompts. Table 4 shows the prompts used for each step discussed in the main paper that employs the LLM (Gemini Pro (Gemini Team Google (2023))). Specifically, we detail the prompts used to: 1) conduct zero-shot image classification with Gemini Pro (Gemini Team Google (2023)), 2) describe a given testing image, and 3) generate class labels descriptions.
The class descriptions were generated using five prompts, as shown in Table 4, with 10 responses generated for each prompt and class, resulting in 50 class descriptions per class label. To encourage diversity in Gemini Pro’s responses, we set the temperature parameter to a high value of 0.99, as done in Pratt et al. (2023). An example of implementing our method, including both classifier construction and the inference process, is shown in Code 1.
Task | Prompt |
Image classification | You are given an image and a list of class labels. Classify the image given the class labels. Answer using a single word if possible. Here are the class labels: {classes} |
Image description | What do you see? Describe any object precisely, including its type or class. |
Class description | 1. Describe what a {class_label} looks like in one or two sentences. 2. How can you identify a {class_label} in one or two sentences? 3. What does a {class_label} look like? Respond with one or two sentences. 4. Describe an image from the internet of a {class_label}. Respond with one or two sentences. 5. A short caption of an image of a {class_label}: |
L/14 | B/16 | B/32 | DistilBERT | RoBERTa | ROUGE-N-F1 | ROUGE-F1 | Ours | |
Top-1 | 52.1 | 52.0 | 55.6 | 43.7 | 28.8 | 54.8 | 54.9 | 70.2 |
In Section 3.1, we visualize examples that highlight the important parts of the inputs contributing to the final predicted class label. In the main paper, we described the approach of sequentially masking out patches from the image and comparing the predicted class with the prediction obtained using the entire unmasked image. Similarly, we follow the same approach for text input by sliding a kernel, masking out words, and comparing the predicted class with our original prediction using inputs without any masking. We used a 2D kernel of size 5050 pixels with a stride of 10 pixels. If there are no highlighted regions in the image due to the small size of the kernel, we enlarge it by 50 until we reach a kernel size of 200200 pixels.
Appendix B Additional Results
In this section, we provide supplementary results to those presented in the main paper. Figure 7 shows the confusion matrix for CLIP (ViT-L/14) (Radford et al. (2021)) and our method across two datasets (Caltech-101 (Fei-Fei et al. (2004)) and CIFAR-100 (Krizhevsky et al. (2009))). The shown results demonstrate that our method enhances classification accuracy and reduces misclassification rates.
@@ Appendix B Additional Results
CLIP (ViT-L/14) (Radford et al. (2021)) | 88.4 | 98.5 | 77.0 | 97.8 | 89.1 | 93.7 | 72.8 | 95.2 | 96.6 | 64.5 |
CuPL (Pratt et al. (2023)) | 91.0 | 98.1 | 79.3 | 98.3 | 92.1 | 94.2 | 77.7 | 99.8 | 96.2 | 68.8 |
@@
CLIP (ViT-L/14) (Radford et al. (2021)) | 0.651 | 0.862 | 0.539 | 0.867 | 0.611 | 0.656 | 0.452 | 0.830 | 0.835 | 0.372 |
CuPL (Pratt et al. (2023)) | 0.665 | 0.851 | 0.573 | 0.889 | 0.652 | 0.637 | 0.480 | 0.902 | 0.795 | 0.395 |
@@
In Table 2, encouraging results were demonstrated by utilizing the feature of initial prediction produced by the LLM (i.e., Gemini Pro (Gemini Team Google (2023))) for zero-shot image classification. Based on these results, one might argue for the direct utilization of Gemini’s class prediction, aiming to match a specific class label from the dataset. However, in several cases, Gemini’s response does not precisely match one of the class labels (as shown in Figure 5). For example, if a ground-truth class label is ‘cat’, Gemini’s response might be ‘The image class is cat’. This discrepancy motivated us to report results of using only Gemini prediction.
In this section, we present additional results from early experiments aimed at utilizing Gemini’s predictions to precisely match one of the class labels in the given dataset. Specifically, we randomly selected 5,000 images from ImageNet (Deng et al. (2009)) for evaluation. While our method, as presented in the main paper, offers a practical way of utilizing Gemini’s predictions, we also present the results of some alternative approaches aimed at precisely identifying one of the dataset class labels, rather than solely relying on the class prediction text generated by Gemini.
Table 5 show the results on the 5,000 images from ImageNet (Deng et al. (2009)) of our main method and alternatives that utilize Gemini’s class prediction to conduct similarity matching with the target dataset class labels. Specifically, we report the results of encoding Gemini’s class prediction using an open-vocabulary language model and measuring the similarity with the encoded class label features. Here, we show the results of using CLIP (ViT-L/14, ViT-B/32, ViT-B/16) (Radford et al. (2021)), DistilBERT (Sanh et al. (2019)), and RoBERTa (Liu et al. (2019)).
In addition, we explored classical text similarity metrics – namely, ROUGE-N-F1 and ROUGE-F1 (Lin (2004)) – rather than encoding both Gemini’s prediction and class labels using an open-vocabulary encoding model. As shown in Table 5, our method, which utilizes Gemini’s class prediction as one of the input features, achieves the best results when compared with the alternative approaches.
In the main paper, we reported the top-1 classification accuracy on several datasets (Deng et al. (2009); Fei-Fei et al. (2004); Bossard et al. (2014); Krause et al. (2013); Krizhevsky et al. (2009); Cimpoi et al. (2014); Zhou et al. (2017); Xiao et al. (2010; 2016); Parkhi et al. (2012)). Table 6 presents the top-5 classification accuracy of our method compared to prior work, while Table 7 shows the Cohen’s kappa coefficient. As can be seen, our method achieves a notable improvement while remaining simple and easy to implement.

### v3

@@ 3.2 Ablation Studies
We conducted a series of ablation studies to explore different versions of our method and investigate the impact of each individual feature, different fusion approaches between the features, and different cross-modal embedding models. Table 3 presents the results of our method using solely the encoded image description, referred to as the description feature (DF). We also report the results obtained by using the encoding of the initial class predictions, termed as the prediction feature (PF), as well as using both DF and PF concurrently as inputs. Additionally, Table 3 shows the results of using image feature (IF) alongside DF or PF as inputs, and finally, we present the results when leveraging all available input features (i.e., DF, PF, and IF).
From the results in Table 3, it is clear that incorporating all three features (i.e., DF, PF, and IF) yields the best performance across most datasets, except for the CIFAR datasets [13]. This discrepancy may arise from the low resolution of CIFAR images (originally 3232 pixels), where utilizing the image feature (IF) may degrade accuracy compared to using only DF and PF.
| CLIP model [21] | |||
@@
| L/14 | B/16 | B/32 | DistilBERT |
| 52.1 | 52.0 | 55.6 | 43.7 |
| RoBERTa | ROUGE-N-F1 | ROUGE-F1 | Ours |
| 28.8 | 54.8 | 54.9 | 70.2 |
Table 4 shows the results of our second set of ablation studies, where we report the results for 5,000 randomly selected images from the ImageNet dataset [5]. We explore the use of different cross-modal embedding models (CLIP [ViT-L/14], CLIP [ViT-B/32], and CLIP [ViT-B/16]) [21], and additionally investigate different fusion approaches. Rather than using the mean feature vector of our input features (DF, PF, and IF), we calculated the similarity between each input feature separately and the dataset class label features. Subsequently, we fused the similarity scores to generate a single similarity score for each class label in each dataset. We explored two fusion methods: averaging and taking the maximum for each class label. As shown, averaging our three features (IF, DF, PF) yields the best results. Generally, averaging is more robust and less sensitive to outliers than taking the maximum, and hence, more representative.
As shown in Tables 2 and 3, encouraging results were demonstrated by utilizing the encoded features from the LLM (i.e., Gemini [9]) for zero-shot image classification. Based on these results, one might argue for the direct utilization of Gemini’s class prediction, aiming to match a specific class label from the dataset. However, in several cases, Gemini’s response does not precisely match one of the class labels. For example, if a ground-truth class label is ‘cat’, Gemini’s response might be ‘The image class is cat’ (see Fig. 1). This discrepancy motivated us to report results of using only Gemini prediction.
Specifically, we present additional results from early experiments aimed at utilizing Gemini’s predictions to precisely match one of the class labels in the given dataset. We randomly selected 5,000 images from ImageNet [5] for evaluation. While our method, offers a practical way of utilizing Gemini’s predictions, we also present the results of some alternative approaches aimed at precisely identifying one of the dataset class labels, rather than solely relying on the class prediction text generated by Gemini.
Table 5 show the results on the 5,000 images from ImageNet [5] of our main method and alternatives that utilize Gemini’s class prediction to conduct similarity matching with the target dataset class labels. Specifically, we report the results of encoding Gemini’s class prediction using an open-vocabulary language model and measuring the similarity with the encoded class label features. Here, we show the results of using CLIP (ViT-L/14, ViT-B/32, ViT-B/16) [21], DistilBERT [22], and RoBERTa [16].
In addition, we explored classical text similarity metrics – namely, ROUGE-N-F1 and ROUGE-F1 [15] – rather than encoding both Gemini’s prediction and class labels using an open-vocabulary encoding model. As shown in Table 5, our method, which utilizes Gemini’s class prediction as one of the input features, achieves the best results when compared with the alternative approaches.
3.3 Computation resources
@@ 4 Limitations
Our method introduces a new approach by leveraging multimodal LLMs to enhance the accuracy of zero-shot image classification. However, it is important to acknowledge that there are still some limitations inherent in our proposed method. Since our method relies on a multimodal LLM to generate the required features (i.e., DF and PF), there may be potential constraints when running on devices with limited computational power, and it may consume more time compared to other methods. Nevertheless, we believe that advancements in LLMs will lead to models that can run efficiently on lower computational power. This would enable broader accessibility and applicability of such models, such as Gemini [9], GPT [3], and LLaMA [23].
@@ 5 Conclusion
We proposed a simple yet effective zero-shot image classification method leveraging multimodal LLMs. Our approach uses an LLM to generate a description of an input image and make an initial class prediction. We then fuse encoded features from the image description, the initial LLM prediction, and the image itself. This straightforward method significantly improves zero-shot classification accuracy compared to existing techniques under the same experimental conditions. Future work will focus on deploying this approach with on-device multimodal LLMs to reduce the computational and communication overhead of cloud-based models.
Supplementary Material
@@ Appendix A Additional Details
In the main paper, we presented our method for zero-shot image classification. Here we give an example of implementing our method, including both classifier construction and the inference process, as shown in Code LABEL:code.
@@
| CLIP (ViT-L/14) [21] | 88.4 | 98.5 | 77.0 | 97.8 | 89.1 | 93.7 | 72.8 | 95.2 | 96.6 | 64.5 |
| CuPL [20] | 91.0 | 98.1 | 79.3 | 98.3 | 92.1 | 94.2 | 77.7 | 99.8 | 96.2 | 68.8 |
@@
| CLIP (ViT-L/14) [21] | 0.651 | 0.862 | 0.539 | 0.867 | 0.611 | 0.656 | 0.452 | 0.830 | 0.835 | 0.372 |
| CuPL [20] | 0.665 | 0.851 | 0.573 | 0.889 | 0.652 | 0.637 | 0.480 | 0.902 | 0.795 | 0.395 |
@@
In the main paper, we visualize examples that highlight the important parts of the inputs contributing to the final predicted class label. We described the approach of sequentially masking out patches from the image and comparing the predicted class with the prediction obtained using the entire unmasked image. Similarly, we follow the same approach for text input by sliding a kernel, masking out words, and comparing the predicted class with our original prediction using inputs without any masking. We used a 2D kernel of size 5050 pixels with a stride of 10 pixels. If there are no highlighted regions in the image due to the small size of the kernel, we enlarge it by 50 until we reach a kernel size of 200200 pixels.
Appendix B Additional Results
In this section, we provide supplementary results to those presented in the main paper. Figure 7 shows the confusion matrix for CLIP (ViT-L/14) [21] and our method across two datasets (Caltech-101 [7] and CIFAR-100 [13]). The shown results demonstrate that our method enhances classification accuracy and reduces misclassification rates.
In the main paper, we reported the top-1 classification accuracy on several datasets [5, 7, 2, 12, 13, 4, 27, 25, 26, 18]. Table 6 presents the top-5 classification accuracy of our method compared to prior work, while Table 7 shows the Cohen’s kappa coefficient. As can be seen, our method achieves a notable improvement while remaining simple and easy to implement.